OpenStudy (anonymous):

You need to administer an IV piggyback of metronidazole 1.5g in 100mL dextrose and water over 60 minutes. You have tubing with a drop factor of 15. How many drops per minute should you infuse?

OpenStudy (popitree):

i think i already solved same type of question for u

OpenStudy (anonymous):

I will post my answer, let me know if i did it right please

OpenStudy (anonymous):

I will post my answer, let me know if i did it right please @popitree

OpenStudy (anonymous):

is it 25 @popitree

OpenStudy (popitree):

it is, 15*100/60 = 150/6=25, so I believe it is correct, if u have access to answers just cross verify to be sure :)

OpenStudy (anonymous):

thanks